OpenStudy (goformit100):

Show that among all quadrilaterals of a given perimeter the square has the largest area.

OpenStudy (goformit100):

@nathan917

OpenStudy (goformit100):

Please help me

OpenStudy (goformit100):

@jim_thompson5910 @him1618

OpenStudy (nathan917):

P = 2l + 2w A = lw => A = l(P - 2l)/2 = lP/2 - l² => dA/dl = P/2 - 2l = 0.....when area is maximum i.e. P = 4l so, 4l = 2l + 2w => 2l = 2w i.e. l = w.......a square I'm not sure.

OpenStudy (goformit100):

Thank you Sir